# Hetero-EUCLID: Interpretable model discovery for heterogeneous hyperelastic materials using stress-unsupervised learning11footnotemark: 1

### v1

@@ 2.2 Segmenting based on displacement data
35:Output: $\mathcal{S}_{k}:k\in{1\dots(i-1)}$}
36:end function
@@
After identifying the inter-segment boundaries, we employ an approach adapted and modified from Shi et al. [45] to register each element in the domain to a material segment. The approach is essentially a single-thread (cannot be parallelized) recursive growth algorithm. It involves choosing a seed element $\mathcal{E}_{\text{seed}}$ from anywhere within the domain ($\mathcal{S}_{\text{domain}}$), with a condition that none of the nodes constituting the element are “flagged” as inter-segment nodes $\left(\boldsymbol{N}_{\text{flag}}\right)$. Next, the seed element and all its neighboring elements $\left(\mathcal{S}_{\text{neighbors }}\right)$ are added into the material segment being discovered $\mathcal{S}_{i}$, with $i$ denoting the index of the material segment. Next, those elements in the neighboring elements set $\mathcal{S}_{\text{neighbors}}$ that were previously absent in $\mathcal{S}_{i}$ and do not contain flagged inter-segment nodes are designated as “growth-frontiers” $\left(\mathcal{S}_{\text{frontier}}\right)$. The neighbors of each element belonging to $\mathcal{S}_{\text{frontier}}$ are then added to the material segment set $\mathcal{S}_{i}$, and some of these neighboring elements are then designated as growth frontiers and the growth continues till a single material segment demarcated by a closed loop of flagged inter-segment nodes is completely discovered. The process then continues with the selection of another seed element from the set $\mathcal{S}_{\text{domain}}\setminus\mathcal{S}_{i}$ until all elements are registered to a material segment set $\mathcal{S}_{i}$. A more formal description of the approach is provided in Algorithm 1, and is depicted in Figure 6.
@@ 2.3 Robustness to noise
To evaluate the robustness of the Hetero-EUCLID framework, we introduce artificial noise into the synthetic surface displacement data obtained from finite element simulations. This was done to mimic experimental displacement measurement noise that would arise due to digital image correlation (DIC) and camera resolution limitations. For the specimen shown in Figure 7, the reference configuration measures $\times 50 mm$ $\times$ 1 mm and is subjected to non-equi-biaxial mechanical loading, with final stretch ratios of $\lambda_{x}=1.6$ and $\lambda_{y}=2.2$. The planar area of the final deformed plate would be $80\times 110$ mm. A typical DIC setup with a 1-megapixel camera would result in an approximate normalized field of view of 0.1 mm/pixel. As reported in Bornert et al. [57], the noise level in DIC setups usually varies between 0.01 to 0.05 pixels, which would correspond to $1\times 10^{-3}$ mm to $5\times 10^{-3}$ mm of noise for the geometry used in the current study. For the current study, we consider two cases of moderate and high noise level by introducing Gaussian noise having zero mean and standard deviation of $\sigma_{u,\text{mod}}=2\times 10^{-3}$ mm and $\sigma_{u,\text{high}}=5\times 10^{-3}$ mm, respectively. Hence, the displacement field is modified as follows
@@
where $(u_{i}^{a})_{\text{FEM}}$ denotes the measured input displacement field without noise and $(u_{i}^{a})_{\text{noise}}$ represents the artificially introduced uncorrelated Gaussian noise at each node $a$ and direction $i$. Next, as would be done with experimental data, the noisy displacement field $\mathcal{U}$ is denoised using Kernel Ridge Regression (KRR) [58] having a Radial Basis Function (RBF) kernel with hyperparameters tuned by random search to minimize the standard deviation of the denoised signal.
@@
    Next, we extract $(f^{a}_{\text{res}})_{\text{het}}$ from $\boldsymbol{b}_{\text{het}}^{\text{residue}}$ as described in Equation 24 and Figure 5(a). We subsample $(f^{a}_{\text{res}})_{\text{het}}$ for each inter-segment boundary and chose 20% of the nodes with the least $(f^{a}_{\text{res}})_{\text{het}}$ for every inter-segment boundary. The use of this scheme for subsampling nodes from the inter-segment boundaries is expected to decrease the number of mislabeled (assigned wrong material segment) nodes contributing towards solving for $\boldsymbol{\theta}$, thereby improving its accuracy and noise-robustness. A detailed study regarding the effect of inter-segment node sub-sampling is provided in B.
The sub-sampling strategy described above significantly reduces the dimensionality of the matrices and vectors involved in Equation 20 thus significantly improving the computational speed. While the expression in Equation 19 remains unchanged, the system defined in Equation 15 is modified to account for the reduced number of degrees of freedom as follows
$\boldsymbol{A}^{\text{fix}}\boldsymbol{\theta}=\boldsymbol{b}^{\text{fix}}\quad\text{with}\quad\boldsymbol{A}^{\text{fix}}\in\mathbb{R}^{3n_{b}\times\mathit{n_{f}n_{c}}},\quad\boldsymbol{b}^{\text{fix}}\in\mathbb{R}^{3n_{b}}$
@@
Equation 28 defines the updated force balance equations incorporating sub-sampled nodal contributions from free, flagged, and fixed regions of the domain.
3 Results and discussion
3.1 Computational data generation
To evaluate the proposed method, we perform finite element simulations using Abaqus/CAE 2025 software to generate a synthetic computational dataset corresponding to various heterogeneity patterns, as illustrated in Figure 7. The model consists of a thin square plate measuring $50\,\text{mm}\times 50\,\text{mm}\times 1\,\text{mm}$, composed of multiple hyperelastic material segments and subjected to quasi-static, displacement-controlled non-equi-biaxial tension. The specimen is analyzed under the assumption of plane stress conditions. The domain is discretized using 6-node linear triangular prism elements (wedge or C3D6 type) as shown in Figure 2. Simulations are conducted under prescribed boundary conditions, and both nodal displacements and boundary-aggregate reaction forces are recorded over a single load step. To emulate realistic DIC measurements, artificial Gaussian white noise is added to the displacement field, as described in Equation 27, with subsequent noise handling procedures detailed in subsection 2.3.
For each heterogeneity pattern shown in Figure 7, we assign different combinations of hyperelastic material models that include a quadratic volumetric strain energy term. These configurations are used to benchmark the performance of the Hetero-EUCLID framework. For the scope of the current work, we use the material library $\boldsymbol{Q}$ with six features as highlighted in Table 1. We employ only $\mathit{n_{f}}=6$ features in the library instead of 24 features in the previous Bayesian-EUCLID work [46]. This is because the number of columns of $\boldsymbol{A}$ matrix would be $\mathit{n_{f}n_{c}}=30$, if there are $n_{c}=5$ material segments in the domain. Using a higher number of features ($n_{f}$) would reduce the speed of computation of the Markov Chain solution and is therefore not attempted in this work. The various hyperelastic material models considered in this work are
  1. A.
    Neo-Hookean model (NH2) [49]
    $W_{\text{NH2}}(\boldsymbol{C};\boldsymbol{\theta})=\theta_{1}(\tilde{I}_{1}-3)+\theta_{6}\left(J-1\right)^{2}$ (29)
  2. B.
    Isihara model (ISH) [48]
    $W_{\text{ISH}}(\boldsymbol{C};\boldsymbol{\theta})=\theta_{1}(\tilde{I}_{1}-3)+\theta_{2}(\tilde{I}_{2}-3)+\theta_{3}(\tilde{I}_{1}-3)^{2}+\theta_{6}\left(J-1\right)^{2}$ (30)
  3. C.
    Haines-Wilson model (HW) [51]
    $W_{\text{HW}}(\boldsymbol{C};\boldsymbol{\theta})=\theta_{1}(\tilde{I}_{1}-3)+\theta_{2}(\tilde{I}_{2}-3)+\theta_{4}(\tilde{I}_{1}-3)(\tilde{I}_{2}-3)+\theta_{5}(\tilde{I}_{1}-3)^{3}+\theta_{6}\left(J-1\right)^{2}$ (31)
Model coefficient Library feature Model coefficient Library feature
$\theta_{1}$ $(\tilde{I}_{1}-3)$ $\theta_{4}$ $(\tilde{I}_{1}-3)(\tilde{I}_{2}-3)$
$\theta_{2}$ $(\tilde{I}_{2}-3)$ $\theta_{5}$ $(\tilde{I}_{1}-3)^{3}$
$\theta_{3}$ $(\tilde{I}_{1}-3)^{2}$ $\theta_{6}$ $(J-1)^{2}$
Table 1: List of features in the library set $\boldsymbol{Q}$ used to represent the Neo-Hookean (29), Isihara (30), and Haines-Wilson (31) constitutive models.
@@ 3.2 Pattern case study
Pattern Type: Cross-shaped inclusion $|$ Mesh elements (C3D6): 10,442 Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{6}$ (MPa) Free nodes sub-sampling Ground Truth, (a) 1 $\text{NH2}_{\text{a}}$ 1.80 - - - - 6.00 - 2 $\text{NH2}_{\text{b}}$ 5.40 - - - - 15.00 No Noise, (f) 1 $\text{NH2}_{\text{a}}$ 1.79 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 6.00 $\pm$ 0.00 2% 2 $\text{NH2}_{\text{b}}$ 5.46 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 15.13 $\pm$ 0.04
Table 2: True vs. predicted material parameters for cross-shaped inclusion pattern without noise using Bayesian-EUCLID framework.
Figure 8: Model discovery for the cross-shaped inclusion pattern (7(a)), with sub-figures (b)–(g) corresponding to the zero noise case. (a) displays the ground truth material segment IDs. (b)–(d) show the displacement fields in the $x$, $y$, and $z$ directions, respectively. (e) presents the residual force norm distribution (Equation 24) from the homogenized model. (f) illustrates the predicted material segments from the segmentation algorithm. (g) highlights the sub-sampled nodes from the $\mathcal{D}^{\text{free}}$ subset used to construct the heterogenized model, which is subsequently solved using the Bayesian framework.
Type 1: Cross-shaped inclusion
7(a) illustrates a heterogeneous specimen with a cross-shaped inclusion embedded in a surrounding matrix material. Through this configuration we aim to examine the effectiveness of the Hetero-EUCLID framework; however, it presents a non-trivial test case due to the presence of sharp corners in inter-segment interfaces. For the cross-shaped inclusion pattern, we begin the analysis for a hyperelastic bi-material setup where both the inclusion and the matrix follow Neo-Hookean constitutive models ($\text{NH2}_{\text{a}}$$\text{NH2}_{\text{b}}$). The domain, modeled as a single-element-thick plate, is discretized using 10,442 uniformly sized wedge-type elements. As depicted in 7(a), the specimen is subjected to quasi-static, displacement-controlled non-equi-biaxial tensile loading over $n_{t}=6$ time steps, reaching final directional stretch ratios of $\lambda_{x}=1.6$ and $\lambda_{y}=2.2$. Only displacement data from the final load-step $n_{t}=6$ is considered for further analysis.
Pattern Type: Cross-shaped inclusion $|$ Mesh elements (C3D6): 10,442 Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{6}$ (MPa) Free nodes sub-sampling Ground Truth, (a) 1 $\text{NH2}_{\text{a}}$ 1.80 - - - - 6.00 - 2 $\text{NH2}_{\text{b}}$ 5.40 - - - - 15.00 Moderate Noise $\sigma_{u}=2\times 10^{-3}$ 1 $\text{NH2}_{\text{a}}$ 1.81 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 6.01 $\pm$ 0.00 2% 2 $\text{NH2}_{\text{b}}$ 5.42 $\pm$ 0.07 0.00 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 14.91 $\pm$ 0.14 High Noise $\sigma_{u}=5\times 10^{-3}$, (f) 2 $\text{NH2}_{\text{a}}$ 1.80 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.99 $\pm$ 0.00 2% 3 $\text{NH2}_{\text{b}}$ 5.25 $\pm$ 0.09 0.00 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 14.40 $\pm$ 0.23
Table 3: True vs. predicted material parameters for cross-shaped inclusion pattern with noise using Bayesian-EUCLID framework.
Figure 9: Model discovery for the cross-shaped inclusion pattern (7(a)), with sub-figures (b)–(g) corresponding to the high noise case. (a) displays the ground truth material segment IDs. (b)–(d) show the displacement fields in the $x$, $y$, and $z$ directions, respectively. (e) presents the residual force norm distribution (Equation 24) from the homogenized model. (f) illustrates the predicted material segments from the segmentation algorithm. (g) highlights the sub-sampled nodes from the $\mathcal{D}^{\text{free}}$ subset used to construct the heterogenized model, which is subsequently solved using the Bayesian framework.

### v2

@@ 2.3 Segmenting based on displacement data
35:  Step 26 ensures that all such elements are assigned to the segment $\mathcal{S}_{0}$.
36:  i1i-1 indicates the number of discovered material segments (excluding $\mathcal{S}_{0}$).
37:  if $\mathcal{S}_{0}\neq\emptyset$ then
38:   for $k=i,(i-1),\dots,1$ do
39:     $\mathcal{S}_{k}\leftarrow\mathcal{S}_{k-1}$
40:   end for
41:   $i\leftarrow i+1$
42:  end if
43:Output: $\mathcal{S}_{k}:k\in{1\dots(i-1)}$} $\triangleright$ $\mathcal{S}_{k}$ is the set of elements contained by the $k^{\text{th}}$ material segment
44:end function
$\mathit{\sigma_{f_{\text{res}}}}=\sqrt{\frac{\sum_{a\in\boldsymbol{{N}_{\text{flag}}}}\left(f^{a}_{\text{res}}-\mu\right)^{2}}{\sum_{a\in\boldsymbol{{N}_{\text{flag}}}}1}}\quad;\quad\mathit{\mu_{f_{\text{res}}}}=\frac{\sum_{a\in\boldsymbol{{N}_{\text{flag}}}}\left(f^{a}_{\text{res}}\right)}{\sum_{a\in\boldsymbol{{N}_{\text{flag}}}}1}$
(a) Type 1: Cross-shaped inclusion
(b) Type 2: Split-domain pattern
(c) Type 3: Multiple inner inclusions
(d) Type 4: Cahn-Hilliard pattern
(e) Type 5: Circular inclusion
(f) Type 6: Circular inclusion with fibers
Figure 7: Graphical illustration of heterogeneous test specimen geometries subjected to non-equi-biaxial tensile loading under three-dimensional plane stress conditions. The heterogeneity patterns include: (a) cross-shaped inclusion, (b) split-domain configuration, (c) multiple non-circular inclusions, (d) Cahn-Hilliard pattern, (e) circular inclusion, and (f) circular inclusion with anisotropic fiber reinforcement.
For zero and moderate noise, it was observed that the average of $f^{a}_{\text{res}}$ (i.e, $\mathit{\mu_{f_{\text{res}}}}$) was significantly smaller than its standard deviation $\mathit{\sigma_{f_{\text{res}}}}$. However, $\mathit{\mu_{f_{\text{res}}}}$ was observed to increase with an increase in displacement noise, while $\mathit{\sigma_{f_{\text{res}}}}$ was nearly constant. This indicates that $\mathit{\mu_{f_{\text{res}}}}$ is a measure of the noise prevalent in the displacement measurement. Further, this implies that $f^{a}_{\text{res}}$ could be high in nodes that are either on the inter-segment boundary or have high displacement noise. In order to disambiguate between the inter-segment nodes and those with high displacement noise, we employ a scalar multiple of $\mathit{\sigma_{f_{\text{res}}}}$ as a lower threshold. In experiments where the displacement noise is sufficiently low (measured by the ratio: $\mathit{\mu_{f_{\text{res}}}}/\mathit{\sigma_{f_{\text{res}}}}$), the inter-segment nodes can be expected to mostly have $f^{a}_{\text{res}}$ above the lower threshold, while the “noisy” nodes would mostly have $f^{a}_{\text{res}}$ below the threshold. This distinction blurs with increasing displacement noise, rendering residual force-based segmentation inaccurate for data with high displacement noise. In subsection 3.6 we provide further discussion on the observed statistical trends for nodal residual force $f^{a}_{\text{res}}$. Notably, Shi et al. [2025] makes similar use of residual forces from a homogenized assumption of their NN surrogate model to highlight the inter-subdomain boundaries, with user-defined lower thresholds.
@@
After identifying the inter-segment boundaries, we employ an approach adapted and modified from Shi et al. [2025] to register each element in the domain to a material segment. The approach is essentially a single-threaded (non-parallelizable) recursive growth algorithm. It involves choosing a seed element $\mathcal{E}_{\text{seed}}$ from anywhere within the domain ($\mathcal{S}_{\text{domain}}$), with a condition that none of the nodes constituting the element are “flagged” as inter-segment nodes $\left(\boldsymbol{N}_{\text{flag}}\right)$. Next, the seed element and all its neighboring elements $\left(\mathcal{S}_{\text{neighbors}}\right)$ are added into the material segment being discovered $\mathcal{S}_{i}$, with $i$ denoting the index of the material segment. Next, those elements in the neighboring elements set $\mathcal{S}_{\text{neighbors}}$ that were previously absent in $\mathcal{S}_{i}$ and do not contain flagged inter-segment nodes are designated as “growth-frontiers” $\left(\mathcal{S}_{\text{frontier}}\right)$. The neighbors of each element belonging to $\mathcal{S}_{\text{frontier}}$ are then added to the material segment set $\mathcal{S}_{i}$, and some of these neighboring elements are then designated as growth frontiers and the growth continues till a single material segment demarcated by a closed loop of flagged inter-segment nodes is completely discovered. The process then continues with the selection of another seed element from the set $\mathcal{S}_{\text{domain}}\setminus\mathcal{S}_{i}$ until all elements are registered to a material segment set $\mathcal{S}_{i}$. A more formal description of the approach is provided in Algorithm 1, and is depicted in Figure 6.
@@
3 Results and discussion
3.1 Computational data generation
To evaluate the proposed method, we perform finite element simulations using Abaqus/CAE 2025 software to generate a synthetic computational dataset corresponding to various heterogeneity patterns, as illustrated in Figure 7. The model consists of a thin square plate measuring $50\,\text{mm}\times 50\,\text{mm}\times 1\,\text{mm}$, composed of multiple hyperelastic material segments and subjected to quasi-static, displacement-controlled non-equi-biaxial tension. The specimen is analyzed under the assumption of plane stress conditions. The domain is discretized using 6-node linear triangular prism elements (wedge or C3D6 type) as shown in Figure 2. Simulations are conducted under prescribed boundary conditions, and both nodal displacements and boundary-aggregate normal reaction forces (not shear forces) are recorded over a single load step. To emulate realistic DIC measurements, artificial Gaussian white noise is added to the displacement field, as described in Equation 34, with subsequent noise handling procedures detailed in subsection 3.2.
For each heterogeneity pattern shown in Figure 7, we assign different combinations of hyperelastic material models that include a quadratic volumetric strain energy term. These configurations are used to benchmark the performance of the Hetero-EUCLID framework. For the scope of the current work, we use the material library $\boldsymbol{Q}$ consisting of nineteen features (Table 1) across two categories of simulation studies. In the first set of studies (Type 1–4), we employ only $\mathit{n_{f}}=6$ features $(\theta_{1}-\theta_{5},\theta_{19})$ from the library $\boldsymbol{Q}$ for the pattern-based case studies corresponding to Type 1–4 configurations (Figure 7a–d), as detailed in subsection 3.3 and 3.4. This is because, for a domain with 5 material segments ($n_{c}=5$), the number of columns of the $\boldsymbol{A}$ matrix would be $\mathit{n_{f}n_{c}}=30$. Incorporating a larger number of features ($\mathit{n_{f}}$) would substantially increase the size of the $\boldsymbol{A}$ matrix, and thereby increase the computational cost of the Markov Chain solution. We have therefore avoided including more than $n_{f}=6$ features in these classes of demonstrative studies, although adding more number of features would not compromise the accuracy of the results [Joshi et al., 2022]. As detailed in subsection 3.5, the second set of simulations evaluates the performance of the proposed framework for the Type 5 and Type 6 configurations (7(e), 7(f)), corresponding to the Ogden and Holzapfel–Gasser–Ogden (HGO) models, respectively. In this study, all features listed in Table 1 are employed while intentionally omitting the Ogden- and HGO-specific terms (refer to Equation 32 and 33) to examine the framework’s robustness in predicting complex hyperelastic material behavior under feature exclusion. The various hyperelastic material models considered in this work are
  1. A.
    Neo-Hookean model (NH2) [Treloar, 1943]
    $W_{\text{NH2}}(\boldsymbol{C};\boldsymbol{\theta})=\theta_{1}(\tilde{I}_{1}-3)+\theta_{19}\left(J-1\right)^{2}$ (29)
  2. B.
    Isihara model (ISH) [Isihara et al., 1951]
    $W_{\text{ISH}}(\boldsymbol{C};\boldsymbol{\theta})=\theta_{1}(\tilde{I}_{1}-3)+\theta_{2}(\tilde{I}_{2}-3)+\theta_{3}(\tilde{I}_{1}-3)^{2}+\theta_{19}\left(J-1\right)^{2}$ (30)
  3. C.
    Haines-Wilson model (HW) [Haines and Wilson, 1979]
    $W_{\text{HW}}(\boldsymbol{C};\boldsymbol{\theta})=\theta_{1}(\tilde{I}_{1}-3)+\theta_{2}(\tilde{I}_{2}-3)+\theta_{4}(\tilde{I}_{1}-3)(\tilde{I}_{2}-3)+\theta_{5}(\tilde{I}_{1}-3)^{3}+\theta_{19}\left(J-1\right)^{2}$ (31)
  4. D.
    Ogden model (OG) [Ogden, 1972]
    $W_{\text{OG}}(\boldsymbol{C};\boldsymbol{\theta})=\frac{2\mu}{\eta}(\tilde{\lambda}_{1}^{\eta}+\tilde{\lambda}_{2}^{\eta}+\tilde{\lambda}_{3}^{\eta}-3)+\theta_{19}\left(J-1\right)^{2}$ (32)
    where $\tilde{\lambda}_{1},\tilde{\lambda}_{2},\tilde{\lambda}_{3}$ are principal stretches with $\tilde{\lambda}_{k}=J^{-1/3}\lambda_{k}$ and $\tilde{\lambda}_{1}\tilde{\lambda}_{2}\tilde{\lambda}_{3}=1$.
  5. E.
    Holzapfel-Gasser-Ogden model (HGO) Holzapfel et al. [2000]
    $W_{\text{HGO}}(\boldsymbol{C};\boldsymbol{\theta})=\theta_{1}(\tilde{I}_{1}-3)+\theta_{18}\left[\frac{(J^{2}-1)}{2}-\ln{(J)}\right]+\frac{k_{1}}{2k_{2}}\left[e^{k_{2}(\tilde{I}_{a}-1)^{2}}-1\right]$ (33)
    where $\tilde{I}_{a}=J^{-2/3}(\boldsymbol{a}\cdot\boldsymbol{Ca})$ denotes anisotropic invariant with one fiber family having orientation angle $\alpha$ with $\boldsymbol{a}=(\cos{\alpha},\sin{\alpha},0)^{T}$.
Model Library Model Library Model Library
coefficient feature coefficient feature coefficient feature
$\theta_{1}$ $(\tilde{I}_{1}-3)$ $\theta_{8}$ $(\tilde{I}_{1}-3)(\tilde{I}_{2}-3)^{2}$ $\theta_{15}$ $(\tilde{I}_{a}-1)^{2}$
$\theta_{2}$ $(\tilde{I}_{2}-3)$ $\theta_{9}$ $(\tilde{I}_{2}-3)^{3}$ $\theta_{16}$ $(\tilde{I}_{a}-1)^{3}$
$\theta_{3}$ $(\tilde{I}_{1}-3)^{2}$ $\theta_{10}$ $(\tilde{I}_{1}-3)^{4}$ $\theta_{17}$ $(\tilde{I}_{a}-1)^{4}$
$\theta_{4}$ $(\tilde{I}_{1}-3)(\tilde{I}_{2}-3)$ $\theta_{11}$ $(\tilde{I}_{1}-3)^{3}(\tilde{I}_{2}-3)$ $\theta_{18}$ $0.5(J^{2}-1)-\ln{(J)}$
$\theta_{5}$ $(\tilde{I}_{2}-3)^{2}$ $\theta_{12}$ $(\tilde{I}_{1}-3)^{2}(\tilde{I}_{2}-3)^{2}$ $\theta_{19}$ $(J-1)^{2}$
$\theta_{6}$ $(\tilde{I}_{1}-3)^{3}$ $\theta_{13}$ $(\tilde{I}_{1}-3)(\tilde{I}_{2}-3)^{3}$ - -
$\theta_{7}$ $(\tilde{I}_{1}-3)^{2}(\tilde{I}_{2}-3)$ $\theta_{14}$ $(\tilde{I}_{2}-3)^{4}$ - -
Table 1: List of features in the library set $\boldsymbol{Q}$ used to represent the Neo-Hookean (29), Isihara (30), Haines-Wilson (31), Ogden (32) and Holzapfel-Gasser-Ogden (33) forms of constitutive models.
@@ 3.2 Robustness to noise
To evaluate the robustness of the Hetero-EUCLID framework, we introduce artificial noise into the synthetic surface displacement data obtained from finite element simulations. This was done to mimic experimental displacement measurement noise that would arise due to digital image correlation (DIC) and camera resolution limitations. For the specimen shown in Figure 7, the reference configuration measures $\times 50 mm$ $\times$ 1 mm and is subjected to non-equi-biaxial mechanical loading, with final stretch ratios of $\lambda_{x}=1.6$ and $\lambda_{y}=2.2$. The planar area of the final deformed plate would be $80\times 110$ mm. A typical DIC setup with a 1-megapixel camera would result in an approximate normalized field of view of 0.1 mm/pixel. As reported in Bornert et al. [2009], the noise level in DIC setups typically varies between 0.01 and 0.05 pixels, which corresponds to $1\times 10^{-3}$ mm and $5\times 10^{-3}$ mm of noise for the geometry used in the current study. For the current study, we consider two cases of moderate and high noise level by introducing Gaussian noise having zero mean and standard deviation of $\sigma_{u,\text{mod}}=2\times 10^{-3}$ mm and $\sigma_{u,\text{high}}=5\times 10^{-3}$ mm, respectively. Hence, the displacement field is modified as follows
@@
where $(u_{i}^{a})_{\text{FEM}}$ denotes the measured input displacement field without noise and $(u_{i}^{a})_{\text{noise}}$ represents the artificially introduced uncorrelated Gaussian noise at each node $a$ and direction $i$. Next, as would be done with experimental data, the noisy displacement field $\mathcal{U}$ is denoised using Kernel Ridge Regression (KRR) [Saunders et al., 1998] having a Radial Basis Function (RBF) kernel with hyperparameters tuned by random search to minimize the standard deviation of the denoised signal. In E we repeat these studies without performing any KRR denoising on the noisy displacement field, while in F we repeat the study for the Type-I geometry (7(a)) while adding spatially correlated displacement noise.
@@
    Next, we extract $(f^{a}_{\text{res}})_{\text{het}}$ from $\boldsymbol{b}_{\text{het}}^{\text{residue}}$ as described in Equation 26 and Figure 5(a). We subsample $(f^{a}_{\text{res}})_{\text{het}}$ for each inter-segment boundary and chose 20% of the nodes with the least $(f^{a}_{\text{res}})_{\text{het}}$ for every inter-segment boundary. The use of this scheme for subsampling nodes from the inter-segment boundaries is expected to decrease the number of mislabeled (i.e., assigned to the wrong material segment) nodes that contribute to solving for $\boldsymbol{\theta}$, thereby improving its accuracy and noise robustness. A detailed study regarding the effect of inter-segment node sub-sampling is provided in D.
The sub-sampling strategy described above significantly reduces the dimensionality of the matrices and vectors involved in Equation 22, thereby substantially improving the computational speed. While the expression in Equation 21 remains unchanged, the system defined in Equation 15 is modified to account for the reduced number of degrees of freedom as follows
$\boldsymbol{A}^{\text{fix}}\boldsymbol{\theta}=\boldsymbol{b}^{\text{fix}}\quad\text{with}\quad\boldsymbol{A}^{\text{fix}}\in\mathbb{R}^{n_{b}\times\mathit{n_{f}n_{c}}},\quad\boldsymbol{b}^{\text{fix}}\in\mathbb{R}^{n_{b}}$
@@
Equation 35 defines the updated force balance equations, incorporating subsampled nodal contributions from the free, flagged, and fixed regions of the domain.
@@ 3.3 Pattern case study
Pattern Type 1: Cross-shaped inclusion
7(a) illustrates a heterogeneous specimen with a cross-shaped inclusion embedded in a surrounding matrix material. Through this configuration, we aim to examine the effectiveness of the Hetero-EUCLID framework; however, it presents a non-trivial test case due to the presence of sharp corners in inter-segment interfaces. For the cross-shaped inclusion pattern, we begin the analysis for a hyperelastic bi-material setup where both the inclusion and the matrix follow Neo-Hookean constitutive models ($\text{NH2}_{\text{a}}$$\text{NH2}_{\text{b}}$). The domain, modeled as a single-element-thick plate, is discretized using 10,442 uniformly sized wedge-type elements. As depicted in 7(a), the specimen is subjected to quasi-static, displacement-controlled non-equi-biaxial tensile loading over $n_{t}=6$ time steps, reaching final directional stretch ratios of $\lambda_{x}=1.6$ and $\lambda_{y}=2.2$. Only displacement data from the final load-step $n_{t}=6$ is considered for further analysis.